. A model that describes the population p(t ) of a fishery in which harvesting takes place at


a constant rate is given by


dp/dt=kp-h



where k and h are positive constants.


(a) Solve the DE subject to P(0) = p


(b) Describe the behavior of the population P(t) for increasing time in the three cases for p=h/k , p=h/k , and 0<p<h/k



(c) Use the results from part (b) to determine whether the fish population will ever go extinct


in finite time, that is, whether there exists a time T>0 such that p(t )=0 . If the population


goes extinct, then find T

1
Expert's answer
2021-11-23T17:15:58-0500

a)

"dP\/dt=kP-h"


"\frac{dP}{kP-h}=dt"


"\frac{1}{k}ln(kP-h)+lnc_1=t"


"c_2(kP-h)=e^{kt}"


"P(t)=\frac{ce^{kt}+h}{k}"


"P(0)=\frac{c+h}{k}=p"


"c=pk-h"


"P(t)=\frac{(pk-h)e^{kt}+h}{k}"


b)

for "p=h\/k" :

"c=0"

"P(t)=\frac{h}{k}"


for "p>h\/k" :

"c>0"

"P(t)>\frac{h}{k}"


for "0<p<h\/k" :

"0<P(t)<\frac{h}{k}"


c)

"P(t)=\frac{(pk-h)e^{kT}+h}{k}=0"


"e^{kT}=\frac{h}{h-pk}"


"T=\frac{ln(h/(h-pk))}{k}"


"T>0" if "h\/(h-pk)>1"